# Is Grokking a Computational Glass Relaxation?

### v1

@@ Abstract
Understanding neural network’s (NN) generalizability remains a central question in deep learning research. The special phenomenon of grokking, where NNs abruptly generalize long after the training performance reaches a near-perfect level, offers a unique window to investigate the underlying mechanisms of NNs’ generalizability. Here we propose an interpretation for grokking by framing it as a computational glass relaxation: viewing NNs as a physical system where parameters are the degrees of freedom and train loss is the system energy, we find memorization process resembles a rapid cooling of liquid into non-equilibrium glassy state at low temperature and the later generalization is like a slow relaxation towards a more stable configuration. This mapping enables us to sample NNs’ Boltzmann entropy (states of density) landscape as a function of training loss and test accuracy. Our experiments in transformers on arithmetic tasks suggests that there is NO entropy barrier in the memorization-to-generalization transition of grokking, challenging previous theory that defines grokking as a first-order phase transition rubin2023grokking . We identify a high-entropy advantage under grokking, an extension of prior work linking entropy to generalizability but much more significant yang2025high . Inspired by grokking’s far-from-equilibrium nature, we develop a toy optimizer WanD based on Wang-landau molecular dynamics, which can eliminate grokking without any constraints and find high-norm generalizing solutions. This provides strictly-defined counterexamples to theory attributing grokking solely to weight norm evolution towards the Goldilocks zone liu2022omnigrok and also suggests new potential ways for optimizer design.
@@ 1 Introduction
Modern deep learning generally relies on heavily overparameterized neural networks (NNs), which presents surprisingly strong generalization capabilities across diverse tasks deng2009imagenet ; vaswani2017attention ; devlin2019bert ; brown2020language . Therefore, understanding the mechanism behind this generalizability is crucial for further advancement in the field. While model generalization performance typically improves with the training performance simultaneously, the grokking phenomenon exhibits dramatically different dynamics, providing an intriguing case for the investigation of generalization. Grokking was first reported in modular arithmetic tasks where researchers used a transformer model power2022grokking . The model can rapidly reach nearly 100% training accuracy but the test accuracy remains very low, suggesting that the model only memorizes the training dataset without actually understanding them. However, after extensive further training, the model exhibits a sudden transition to a state of high generalization where the test accuracy is (almost) at perfect level. This significantly delayed generalization distinguishes grokking from the typical learning processes and raises fundamental challenges towards the existing generalization theory power2022grokking ; zhang2021understanding ; kumar2023grokking .
@@
Previous works have shown that the loss landscape of overparameterized NNs possesses infinitely many minima; they exhibit different generalizability and are connected with each other to form a single low-loss manifold draxler2018essentially . Therefore, many studies have attempted to propose geometric indicators that correlate with the generalizability of minima. One representative hypothesis is that wider minima generalizes better than their narrower analogs hochreiter1997flat . However, it is difficult to measure or even rigorously define the width of a minimum in such a high-dimensional space. While some indicators like sharpness keskar2016large correlates with model generalization empirically, they have also been questioned for the validity in measuring generalizability alone, due to the general scaling invariance in NNs dinh2017sharp . Other indicators such as the Hessian matrix are too expensive to calculate in deep learning applications keskar2016large . These underscore the limitations of using local geometry alone to predict generalization and thus hinder the understanding of complex generalization dynamics like grokking.
@@
Overcoming the limitations of purely geometric descriptions, a very recent work has explored applying concepts from statistical mechanics to NNs, which conceptualizes NN as a physical system where parameters are the degrees of freedom and uses Boltzmann entropy calculated over parameter space to characterize the loss landscape yang2025high . In this context, entropy measures the logarithm of the volume of solutions in the parameter space. The work reveals a high-entropy advantage for NNs’ generalizability, where the high entropy solutions (the thermodynamic equilibrium state in the context of physics) tend to exhibit better generalization than solutions obtained via classical optimizers like SGD or Adam yang2025high . This motivates us to employ the entropy framework to investigate the dynamics of grokking. We hypothesize that the transition from memorization (pre-grokking) to generalization (grokking) might involve the network navigating towards the higher-entropy regions. Following the reported protocol yang2025high , we employed the Wang-Landau Molecular Dynamics (WLMD) algorithm to sample transformer neural networks’ entropy $S(L_{train},A_{test})$ as a function of the training loss $L_{train}$ and the test accuracy $A_{test}$, for three different modular arithmetic tasks with varying complexity. We find that the grokking process can be mapped to a computational glass relaxation: the training loss of a NN is equivalent to the internal energy of a glass, the memorization training mimics a rapid liquid cooling to form a non-equilibrium glass at low temperature, and the eventual grokking emerges as a slow relaxation process towards a more stable, higher-entropy configuration that generalizes.
@@
The paper is organized as follows: In Section 2, we first review the background of generalization, grokking, Boltzmann entropy, and glass physics. In Section 3, we introduce in detail how we sample the entropy landscape of an NN with WLMD. In Section 4, we first present results from three different modular arithmetic tasks, where two of them exhibits grokking while the other one is unlearnable. The constructed entropy landscapes indicate that there is no entropy barrier between the memorization and generalization states. Therefore, grokking is NOT a first-order phase transition, which contrasts with Rubin et al.’s previous work rubin2023grokking . A much more significant high-entropy advantage has been identified for the grokking NNs, which we referred as high-entropy advantage under grokking in this work. We then eliminate grokking by training with constrained weight norm liu2022omnigrok and find the high-entropy advantage is significantly reduced. These findings inspire us to develop a toy optimizer, WanD, presented in Section 5. We demonstrate WanD can find generalized solutions in a competitive efficiency to AdamW while almost eliminating grokking, without any weight norm constraint or output modifications. Finally, in Section 6 we further discuss the connection between our work and previous studies in grokking, and summarize how existing results link grokking to glass relaxation processes. We also discuss directions for future research.
@@
  • Grokking is NOT a first-order phase transition. By conceptualizing neural networks as a physical system and sampling its entropy landscape, we find there is no entropy barrier between the pre-grokking memorization state and the final generalization state, arguing against the first-order phase transition interpretation rubin2023grokking . Instead, grokking has similar behavior to the formation and relaxation of glass, where the pre-grokking memorization states can be attributed to the fast convergence of train loss minimization, resembles the rapid cooling of a liquid during glass transition. The subsequent grokking process can therefore be recognized as a slow, computational glass relaxation towards a more stable, generalizing configuration. Such an analogy can describe many reported observations for grokking and inspire potential research directions for NNs, as we detailed later.
  • The high-entropy advantage under grokking is a unique property of generalization. We found that in the entropy landscape of the grokking NNs, there is a huge gap between the training curve and the max-entropy curve, which constitutes the previously reported high-entropy advantage for NNs’ generalizability yang2025high . We term it the high-entropy advantage under grokking. Furthermore, we show that after eliminating grokking by controlling the weight norm, the high-entropy advantage is much less prominent but still exists. This confirms that the NN without grokking does not have glass behavior, as they are much closer to the equilibrium state, supporting the correspondence between grokking and glass relaxation. It also shows that there is a broad correlation between the high-entropy solutions and the generalizability of NNs.
  • We design a physics-inspired toy optimizer, WanD, which finds high-norm generalizing solutions without grokking. WanD is an optimizer based on the Wang-Landau molecular dynamics method. In modular addition task training, WanD has similar time efficiency and generalization performance as AdamW but can eliminate grokking without any explicit regularization (like weight norm constraints). Notably, WanD often finds high-norm generalizing solutions, which challenges theory attributing grokking solely to weight decay liu2022omnigrok . This also agrees with the conclusion of Kumar et al. kumar2023grokking that regularization is not always necessary for grokking, but in a strictly-defined situation without output scaling. We argue that grokking is an issue caused by optimization dynamics and can be solved by just improving the optimizer. We believe that WanD points out a potential direction to design new optimizers which has an advantage in finding higher entropy solutions with better generalizability, and our research can encourage further work on applying well-established statistical-physical methodologies for deep learning research.
@@ 2 Related Work
Generalization has long been linked to flat minima in the loss landscape hochreiter1997flat . In an over-parameterized NN, these minima form a connected low-loss manifold draxler2018essentially ; shevchenko2020landscape . Some optimizers like SGD are believed to prefer flatter minima due to its inherent gradient noise xie2020diffusion , explaining why SGD can usually yield better generalization than other optimizers xie2020diffusion . Therefore, many works focus on developing flatness-based measures and observe correlations with generalizability empirically keskar2016large ; dziugaite2017computing ; jiang2019fantastic ; foret2020sharpness . However, the usage of flatness as a simple predictor of generalization has been challenged by many works, mainly because it has the dimension of weight feng2023activity . Dinh et al. dinh2017sharp changed the loss landscape by reconstructing the fitting function without affecting the generalizability, and therefore demonstrated that sharp minima can also generalize. Neyshabur et al. neyshabur2017exploring and Feng et al. feng2023activity discussed that flatness itself is not sufficient to ensure generalization, but needs to be combined with the weight norm to obtain a suitable complexity measure. Our results support this because the combination of flatness and weight norm actually reflects the volume of the set of solutions with the same generalizability in the parameter space, as we will explain in Section 6.
@@
Grokking is an anomalous generalization phenomenon that was first discovered by Power et al. power2022grokking in arithmetic tasks, describing a model’s sudden transition from random-guess to (nearly) perfect test accuracy long after achieving perfect training performance. Different mechanisms have been proposed to explain this delayed generalization, including: (a) shifts from poor to good representation learning kumar2023grokking ; varma2023explaining ; kunin2024get ; (b) transitions from dense to sparse predictive subnetworks merrill2023tale ; (c) slingshot mechanism for adaptive optimizers thilak2022slingshot ; (d) first-order phase transition in the network’s internal representation rubin2023grokking ; and (e) evolution of weight norms towards the Goldilocks zone liu2022omnigrok . Our analysis from the entropy perspective can explain the observations in (a)(b)(c), but challenges (d)(e) with well-defined counterexamples. Detailed discussion on connections with existing work in grokking are presented in Section 6.
@@
Boltzmann Entropy defines the number of microstates for a given macrostate in a thermodynamic system. For example, entropy can be defined by the volume in phase space with a certain energy value: $S(U_{0})=k_{B}lnV|_{U=U_{0}}$, where $k_{B}$ is Boltzmann constant and $V$ represents the volume in phase space that correspond to the internal energy $U_{0}$. Yang et al. yang2025high applied the concept of Boltzmann entropy to the NNs. This allowed them to discarded the local perspective of sharpness and weight norm and established a mapping relationship between generalization and the volume of parameter space. They found that high-entropy NN states usually generalize better than states trained with classical optimizers like SGD, which they define as the high-entropy advantage. This inspired us to use the same approach to study the relationship between entropy and generalization in grokking tasks.
@@
Glass state is a physical state of matter, and is usually formed by rapidly cooling or compressing a liquid. The glass state is not in thermodynamic equilibrium. Therefore, at appropriate temperatures, the glass state will undergo structural relaxation and gradually transform into a better equilibrated state through a kinetic process riechers2022predicting . Winter and Janssen winter2025glassy compared deep NNs with structural glass and found many similarities, including relaxation dynamics at low temperatures. As we will show in this paper, grokking is thermodynamically similar to glass relaxation, so a promising future direction is to explore whether the conclusions in winter2025glassy holds in our setup.
@@ 3 Methodology
In this section, we will first give a definition of the modular arithmetic task, then introduce the concept of entropy into NNs, and finally introduce our sampling method: Wang-Landau molecular dynamics algorithm.
@@ 3.1 Problem definition
The training part of our experiment is same as the setup of Nanda et al. nanda2023progress , that is, a one-layer transformer model trained in modular arithmetic tasks. The data consists of two parts: prompts $\langle x\rangle\langle y\rangle\langle p\rangle$ and answer $\langle x\circ y\rangle$. In the three tasks we study, the answers are determined by the following equations:
[…]
@@ 3.2 Entropy in neural networks
There are many ways to define entropy in physics and information theory. In this work, we apply the definition of Boltzmann entropy $S=k_{B}\ln W$ to NNs. Here, $k_{B}$ is the Boltzmann constant, which we set to 1 for simplicity. In physics, $W$ is the volume of the phase space that corresponds to a particular set of observables (energy, magnetization, etc.). For an NN, we define $W$ as the volume of the parameter space that corresponds to a particular training loss and test accuracy, and therefore $S$ is also a function of these two quantities. To ensure that $W$ is finite, we limit our study to a hypercubic region of the parameter space by constraining the NN parameters to the range $[-\frac{3}{\sqrt{m}},\frac{3}{\sqrt{m}}]$ in Sec. 4.1, where $m$ is the number of neurons in that layer. In Sec. 4.2 we limit our study to a hyperspherical shell of weight norm 30. Considering grokking often occurs when the training loss is small, our results are presented in terms of its logarithm, $S[\ln(L_{train}),A_{test}]$.
The entropy landscape we calculates cannot reach infinite resolution. For computational purposes, we divide $S[\ln(L_{train}),A_{test}]$, which is a 2D function, into rectangular bins, where each bin corresponds to the entropy of the NN parameters within a certain range of training loss and test accuracy. The Wang-Landau Molecular Dynamics method allows us to calculate $S[\ln(L_{train}),A_{test}]$ with an arbitrary zero point. Thus, the absolute value of the entropy is meaningless, but the entropy difference between any two bins are meaningful. Therefore, we can subtract a constant from the entropy in each bin so that the minima-bin entropy is always 0.

### v2

@@ Abstract
Understanding neural network’s (NN) generalizability remains a central question in deep learning research. The special phenomenon of grokking, where NNs abruptly generalize long after the training performance reaches a near-perfect level, offers a unique window to investigate the underlying mechanisms of NNs’ generalizability. Here we propose an interpretation for grokking by framing it as a computational glass relaxation: viewing NNs as a physical system where parameters are the degrees of freedom and train loss is the system energy, we find memorization process resembles a rapid cooling of liquid into non-equilibrium glassy state at low temperature and the later generalization is like a slow relaxation towards a more stable configuration. This mapping enables us to sample NNs’ Boltzmann entropy (density of states) landscape as a function of training loss and test accuracy. Our experiments in transformers on arithmetic tasks suggests that there is NO entropy barrier in the memorization-to-generalization transition of grokking, challenging previous theory that defines grokking as a first-order phase transition rubin2023grokking . We identify a high-entropy advantage under grokking, an extension of prior work linking entropy to generalizability but much more significant yang2025high . Inspired by grokking’s far-from-equilibrium nature, we develop a toy optimizer, WanD, based on Wang-Landau Molecular Dynamics, which can eliminate grokking without any constraints and find high-norm generalizing solutions. This provides strictly-defined counterexamples to theory attributing grokking solely to weight norm evolution towards the Goldilocks zone liu2022omnigrok and also suggests new potential ways for optimizer design.
@@ 1 Introduction
Modern deep learning generally relies on heavily overparameterized neural networks (NNs), which presents surprisingly strong generalization capabilities across diverse tasks deng2009imagenet ; vaswani2017attention ; devlin2019bert ; brown2020language . Therefore, understanding the mechanism behind this generalizability is crucial for further advancement in the field. While model generalization performance typically improves with the training performance simultaneously, the grokking phenomenon exhibits dramatically different dynamics, providing an intriguing case for the investigation of generalization. Grokking was first reported in modular arithmetic tasks where researchers used a transformer model power2022grokking . The model can rapidly reach nearly 100% training accuracy but the test accuracy remains very low, suggesting that the model only memorizes the training dataset without actually understanding them. However, after extensive further training, the model exhibits a sudden transition to a state of high generalization where the test accuracy is (almost) at perfect level. This significantly delayed generalization distinguishes grokking from typical learning processes and raises fundamental challenges towards the existing generalization theory power2022grokking ; zhang2021understanding ; kumar2023grokking .
@@
Previous works have shown that the loss landscape of overparameterized NNs possesses infinitely many minima; they exhibit different generalizability and are connected with each other to form a single low-loss manifold draxler2018essentially . Therefore, many studies have attempted to propose geometric indicators that correlate with the generalizability of minima. One representative hypothesis is that wider minima generalizes better than their narrower analogs hochreiter1997flat . However, it is difficult to measure or even rigorously define the width of a minimum in such a high-dimensional space. While some indicators like sharpness keskar2016large correlates with model generalization empirically, they have also been questioned for the validity in measuring generalizability alone, due to the general scaling invariance in NNs dinh2017sharp . Other indicators such as the Hessian matrix are too expensive to calculate in deep learning applications keskar2016large . These underscore the limitations of using local geometry alone to predict generalization, which hinders the understanding of complex generalization dynamics like grokking.
@@
Overcoming the limitations of purely geometric descriptions, a very recent work has explored applying concepts from statistical mechanics to NNs, which conceptualizes NN as a physical system where parameters are the degrees of freedom and uses Boltzmann entropy calculated over parameter space to characterize the loss landscape yang2025high . In this context, entropy measures the logarithm of the volume of solutions in the parameter space. The work reveals a high-entropy advantage for NNs’ generalizability, where the high entropy solutions (the thermodynamic equilibrium state in the context of physics) tend to exhibit better generalization than solutions obtained via classical optimizers like SGD or Adam yang2025high . This motivates us to employ the entropy framework to investigate the dynamics of grokking. We hypothesize that the transition from memorization (pre-grokking) to generalization (grokking) might involve the network navigating towards the higher-entropy regions. Following the reported protocol yang2025high , we employed the Wang-Landau Molecular Dynamics (WLMD) algorithm to sample transformer neural networks’ entropy $S(ln(L_{train}),A_{test})$ as a function of the logarithm of training loss $ln(L_{train})$ and the test accuracy $A_{test}$, for three different modular arithmetic tasks with varying complexity. We find that the grokking process can be mapped to a computational glass relaxation: the training loss of a NN is equivalent to the internal energy of a glass, the memorization training mimics a rapid liquid cooling to form a non-equilibrium glass at low temperature, and the eventual grokking emerges as a slow relaxation process towards a more stable, higher-entropy configuration that generalizes.
@@
The paper is organized as follows: In Section 2, we review the background of generalization, grokking, Boltzmann entropy, and glass physics. In Section 3, we introduce in detail how we sample the entropy landscape of an NN with WLMD. In Section 4, we first present entropy landscapes from three different modular arithmetic tasks, where two of them exhibits grokking while the other one is unlearnable. The results indicate that there is no entropy barrier between the memorization and generalization states. Therefore, grokking is NOT a first-order phase transition, which contrasts with Rubin et al.’s previous work rubin2023grokking . The grokking NNs also exhibit much more significant high-entropy advantages than those studied in Ref. yang2025high . We then eliminate grokking by training with constrained weight norm liu2022omnigrok and find that the high-entropy advantage is significantly reduced. These findings inspire us to develop a toy optimizer, WanD, presented in Section 5. We demonstrate that WanD can find generalizing solutions in a competitive efficiency to AdamW while almost eliminating grokking, without any weight-norm constraint or output modifications. Finally, in Section 6 we further discuss related work and situate previous findings on grokking within our proposed framework of computational glass relaxation. We also discuss directions for future research.
@@
  • Grokking is NOT a first-order phase transition. By conceptualizing neural networks as a physical system and sampling its entropy landscape, we find there is no entropy barrier between the pre-grokking memorization state and the final generalizing state, arguing against the first-order phase transition interpretation rubin2023grokking . Instead, grokking has similar behavior to the formation and relaxation of glass, where the pre-grokking memorization states can be attributed to the fast convergence of train loss minimization, and resembles the rapid cooling of a liquid during glass transition. The subsequent grokking process can therefore be recognized as a slow, computational glass relaxation towards a more stable, generalizing configuration. Such an analogy can describe many reported observations for grokking and inspire potential research directions for NNs, as we detailed later.
  • The high-entropy advantage under grokking is a unique property of generalization. We find that in the entropy landscape of the grokking NNs, there is a huge gap between the training curve and the max-entropy curve, which was previously reported as “high-entropy advantage in generalizability” for other NNs and datasets yang2025high . The high-entropy advantage under grokking, as we found here, is much larger. Furthermore, we show that after eliminating grokking by controlling the weight norm, the high-entropy advantage is much less prominent but still exists. This confirms that the NN without grokking does not have glassy behavior, as they are much closer to the equilibrium state, supporting the correspondence between grokking and glass relaxation. It also reinforces the understanding that higher entropy correlates with better NNs’ generalizability yang2025high , broadening its empirical support to our distinct model and task.
  • We design a physics-inspired toy optimizer, WanD, which finds high-norm generalizing solutions without grokking. WanD is an optimizer based on the Wang-Landau molecular dynamics method. In the modular addition task, WanD has similar time efficiency and generalization performance as AdamW but can eliminate grokking without any explicit regularization (like weight norm constraints). Notably, WanD often finds high-norm generalizing solutions, which challenges a theory attributing grokking solely to weight decay liu2022omnigrok . This also agrees with the conclusion of Kumar et al. kumar2023grokking that regularization is not always necessary for grokking, but in a strictly-defined situation without output scaling. We argue that grokking is an issue caused by optimization dynamics and can be solved by just improving the optimizer. We believe that WanD points out a potential direction to design new optimizers which has an advantage in finding higher entropy solutions with better generalizability, and our research can encourage further work on applying well-established statistical-physical methodologies for deep learning research.
@@ 2 Related Work
Generalization has long been linked to flat minima in the loss landscape hochreiter1997flat . In an over-parameterized NN, different minima form a connected low-loss manifold draxler2018essentially ; shevchenko2020landscape . Some optimizers like SGD are believed to prefer flatter minima due to its inherent gradient noise xie2020diffusion , explaining why SGD can usually yield better generalization than other optimizers xie2020diffusion . Therefore, many works focus on developing flatness-based measures and observe correlations with generalizability empirically keskar2016large ; dziugaite2017computing ; jiang2019fantastic ; foret2020sharpness . However, the usage of flatness as a simple predictor of generalization has been challenged by many works, mainly because it has the dimension of weight feng2023activity . Dinh et al. dinh2017sharp changed the loss landscape by reconstructing the fitting function without affecting the generalizability, and therefore demonstrated that sharp minima can also generalize. Neyshabur et al. neyshabur2017exploring and Feng et al. feng2023activity discussed that flatness itself is not sufficient to ensure generalization, but needs to be combined with the weight norm to obtain a suitable complexity measure. Our results support this because the combination of flatness and weight norm actually reflects the entropy, the logarithm of the parameter space volume, of the solutions with the same generalizability.
@@
Grokking is an anomalous generalization phenomenon that was first discovered by Power et al. power2022grokking in arithmetic tasks, describing a model’s sudden transition from random-guess to (nearly) perfect test accuracy long after achieving perfect training performance. Different mechanisms have been proposed to explain this delayed generalization, including: (a) shifts from poor to good representation learning kumar2023grokking ; varma2023explaining ; kunin2024get ; liu2022towards ; (b) transitions from dense to sparse predictive subnetworks merrill2023tale ; (c) slingshot mechanism for adaptive optimizers thilak2022slingshot ; (d) first-order phase transition in the network’s internal representation rubin2023grokking ; and (e) evolution of weight norms towards the Goldilocks zone liu2022omnigrok . Our analysis from the entropy perspective is consistent with (a)(b)(c), but challenges (d)(e) with well-defined counterexamples. Detailed discussion on connections with existing work in grokking are presented in Section 6.
@@
Boltzmann Entropy defines the number of microstates for a given macrostate in a thermodynamic system. For example, entropy can be defined by the volume in phase space with a certain energy value: $S(U_{0})=k_{B}lnV|_{U=U_{0}}$, where $k_{B}$ is Boltzmann constant and $V$ represents the volume in phase space that correspond to the internal energy $U_{0}$. Yang et al. yang2025high applied the concept of Boltzmann entropy to the NNs. This allowed them to improve upon the local perspective of sharpness and weight norm, and established a mapping relationship between generalization and the volume of parameter space. They found that high-entropy NN states usually generalize better than states trained with classical optimizers like SGD, which they define as the high-entropy advantage. This inspired us to use the same approach to study the relationship between entropy and generalization in grokking tasks.
@@
Glass state is a physical state of matter, and is usually formed by rapidly cooling or compressing a liquid. The glass state is not in thermodynamic equilibrium. Therefore, at appropriate temperatures, the glass state will undergo structural relaxation- a process where the atomic or molecular arrangement in a non-equilibrium glass gradually evolves toward a more stable, more equilibrated configuration over time. This relaxation process is generally regarded as a kinetic process riechers2022predicting . Winter and Janssen winter2025glassy compared deep NNs with structural glass and found many similarities, including the relaxation dynamics at low temperatures. As we will show in this paper, grokking is thermodynamically similar to glass relaxation, so a promising future direction is to explore whether the conclusions in winter2025glassy holds in our setup.
@@ 3 Methodology
In this section, we will first define the modular arithmetic task, then introduce the concept of entropy into NNs, and finally present our sampling method: the Wang-Landau Molecular Dynamics algorithm.
@@ 3.1 Problem definition
The training part of our experiment is same as the setup of Nanda et al. nanda2023progress , that is, a one-layer transformer model trained in modular arithmetic tasks. The data consists of two parts: prompts $\langle x\rangle\langle y\rangle\langle p\rangle\$and answer $\langle x\circ y\rangle$. In the three tasks we study, the answers are determined by the following equations:
$x\circ y=x+y\ mod\ p\ \mathrm{for}0\leq x,y<p$ (1)
$x\circ y=x^{2}+y\ mod\ p\ \mathrm{for}0\leq x,y<p$
$x\circ y=x^{3}+xy^{2}+y\ mod\ p\ \mathrm{for}0\leq x,y<p$
@@ 3.2 Entropy in neural networks
There are many ways to define entropy in physics and information theory. In this work, we apply the definition of Boltzmann entropy $S=k_{B}\ln V$ to NNs. Here, $k_{B}$ is the Boltzmann constant, which we set to 1 for simplicity. In physics, $V$ is the volume of the phase space that corresponds to a particular set of observables (energy, magnetization, etc.). For an NN, we define $V$ as the volume of the parameter space that corresponds to a particular training loss and test accuracy, and therefore $S$ is also a function of these two quantities. To ensure that $V$ is finite, we limit the parameter space by setting different reflective bounds for the parameters of different layers. These bounds are designed to cover the range of parameters typically reached during normal training while avoiding being excessively large. In Sec. 4.2, we avoid applying these reflective bounds to study the entropy advantage when grokking is suppressed. Instead, we limit our study to a hyperspherical shell of weight norm 30, a method known to effectively prevent grokking liu2022omnigrok . Considering grokking often occurs when the training loss is small, our results are presented in terms of its logarithm, $S(ln(L_{train}),A_{test})$. Here, $L_{train}$ denotes the training cross-entropy loss, and $A_{test}$ represents the smoothed test accuracy.
The entropy landscape we calculate cannot reach infinite resolution. For computational purposes, we divide $S(ln(L_{train}),A_{test})$, which is a 2D function, into rectangular bins, where each bin corresponds to the entropy of the NN parameters within a certain range of training loss and test accuracy. The Wang-Landau Molecular Dynamics method allows us to calculate $S(ln(L_{train}),A_{test})$ with an arbitrary zero point. In other words, the absolute value of the entropy is meaningless, but the entropy difference between any two bins is meaningful. Therefore, we can subtract a constant from the entropy in each bin so that the minima-bin entropy is always 0.